What is the mass of silver that can be prepared from 3.00 g of copper metal?
1
Expert's answer
2017-11-22T13:45:06-0500
Silver reduction takes place in the following reaction:
Cu + 2Ag+ -> Cu2+ + 2Ag
So each mole of Cu releases 2 moles of Ag.
In other words 64g of copper release 216 g of silver.
Let’s make a proportion:
64 g Cu ---- 216 g Ag
3 g Cu ---- X g Ag
X = (3×216)/64 = 10.125 (g)
Answer
10,125 g of silver that can be prepared from 3.00 g of copper metal.